The intensity of sound from a point source is 1.0 × 10-8 W m-2 at a distance of 5.0 m from the source. What will be the intensity at a distance of 25m from the source?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Here I base 1 = 1.0 × 10^–8 W base 1/m^2 ; I base2 = ? r base 1 = 5.0 m, r base 2 = 25 m. We know that I ∝ 1/r^2 ⇒l base 1 r base 1^2 ⇒ l base 2 = l base 1 r^2 base 1/r^2 base 2 = 1.0 * 10^-8 * 25/625 = 4.0 * 10^-10 W/m^2. Here I base 1 = 1.0 × 10^–8 W base 1/m^2 ; I base2 = ? r base 1 = 5.0 m, r base 2 = 25 m. We know that I ∝ 1/r^2 ⇒l base 1 r base 1^2 ⇒ l base 2 = l base 1 r^2 base 1/r^2 base 2 = 1.0 * 10^-8 * 25/625 = 4.0 * 10^-10 W/m^2